Solve for x using logarithms:
Solve for x: Log base 4 of(x-6)+ log base 4 of(x+2)=2logx

Answer by stanbon(75887)

Solve for x: Log base 4 of(x-6)+ log base 4 of(x+2)=2logx
----
log4[(x-6)(x+2)] = log4(x^2)
----
x^2-4x-12 = x^2
----
-4x-12 = 0
4x = -12
x = -3
-----
But x cannot be -3 because the
original equation has a terms "log4(x+2)"
and 2log4(x). These would be
meaningless if x = -3
=================
Final answer: No solution.
=================
Cheers,
Stan H.